OpenStudy (anonymous):

I am working on this problem This year, a 30-second commercial to be shown during the Super Bowl will cost $4 million. How much would a 20-second commercial cost? A 40-second commercial? An estimated 110 million people will be watching this year. About how much money is being spent per person in a 30-second ad? So far this is what I have with help fromlucaz here but I am not sure I am doing it right. This is what I have so far am I correct? set up as a proportion and cross multiply 4000000/30=X/20 then I got 80000000 by multiplying 4000000 by 20 then I divided 80000000 by 30 =2666666.667

OpenStudy (anonymous):

The first question: Divide 30 by $4 million. This gives you the cost per second and then multiply by 20 and 40 to answer the first two questions.

OpenStudy (anonymous):

Sorry. I meant divide 4 million by 30.

OpenStudy (phi):

yes, you are doing it correctly. But I would work in millions of dollars (to avoid all those zeros) in other words do \[ \frac{4}{30}= \frac{x}{20}\] and you get \[ x= \frac{4 \cdot 20}{30} = \frac{8}{3}= 2 \frac{2}{3} \text{ million dollars}\]

OpenStudy (anonymous):

IThanks I am in 6th grade and these problems come with no instructions. Thanks to both of you. for the cost person I am dividing 4M by 110M and getting 0..36... rounding to about 4 cents a person is this correct?

OpenStudy (31356):

Phi, is correct.

OpenStudy (phi):

4/110 = 3.64 cents per person ( I would not round it to 4 cents)

OpenStudy (anonymous):

Thanks so much I wanted someone to check my work because I was unsure. Now I can practice some problems like this to remember. This is my weekend word problem from school.

OpenStudy (31356):

Great idea.

OpenStudy (anonymous):

I keep losing the connection thanks